### v1

# Serving Our AI Overlords: Redesigning Data Systems for Agents

Abstract.
Large Language Model (LLM) agents, acting on their users’ behalf to manipulate and analyze data, are likely to become the dominant workload for data systems in the future. When working with data, agents employ a high-throughput process of exploration and solution formulation for the given task, one we call agentic speculation. The sheer volume and inefficiencies of agentic speculation can pose challenges for present-day data systems. We argue that data systems need to adapt to more natively support agentic workloads. We take advantage of the characteristics of agentic speculation that we identify, i.e., scale, heterogeneity, redundancy, and steerability—to outline a number of new research opportunities for a new agent-first data systems architecture, ranging from new query interfaces, to new query processing techniques, to new agentic memory stores.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
Powered by Large Language Models (LLMs) that can reason, invoke tools, author code, and communicate with each other, we are on the precipice of a new agentic revolution that will transform how data systems are used. Modern LLMs are far more efficient internally, matching the capabilities of those orders of magnitude larger just a year ago, and growing ever more effective at understanding and manipulating both structured and unstructured data. As they become both cheap and capable, future LLM agents will act on users’ behalf: extracting, analyzing, transforming, and updating data—potentially becoming the dominant workload for data systems.
While LLM agents may match human reasoning capabilities, they won’t possess grounding—an awareness of the underlying data and characteristics of the data systems on which the data is stored. However, they can make up for this lack of grounding by tirelessly working through possible solutions to a given data transformation task, far more than any human could or would. Each individual LLM agent can theoretically issue hundreds or thousands of requests a second [^1] https://developer.nvidia.com/deep-learning-performance-training-inference/ai-inference, with this rate scaling with the number of LLM agents. Many of these requests are not attempts at a solution, but are instead part of an exploratory process of metadata discovery (e.g., table schemas, column statistics), coupled with partial solutions and validation. We refer to this combination of discovery and solution formulation as agentic speculation—i.e., high-throughput, exploratory querying to identify the best course of action.
Agentic speculation represents a substantial departure from present-day data systems workloads, which are either more intermittent (e.g., from humans or tools operating on their behalf) or more targeted (e.g., from end-user applications). Consider an army of LLM agents tasked with finding reasons for why profits in coffee bean sales in Berkeley was low this year relative to last. Since they are not limited by human cognitive bandwidth and response times, an army of agents could employ an enormous volume of queries to data systems, far more than any human could—all for a single task. Many of these queries are likely wasteful, and are simply providing the agents grounding. As another example, if an LLM agent is tasked with identifying a new crew for a delayed flight, it would need to consider various hypothetical transactions to surface to a human decision maker, each with dozens of updates to various databases. [^2] Example thanks to Keshav Murthy at Couchbase. For such tasks, agents may explore many alternatives in parallel by forking database state, running speculative updates, and rolling back branches. Overall, as agentic workloads become more and more prevalent, the sheer scale and inefficiencies of agentic speculation will become the bottleneck, and our data systems will need to evolve in response.
So we ask the question: how can data systems evolve to better support agentic workloads? In particular, can data systems natively—and efficiently—support agentic speculation, helping LLM agents determine the best course of action? This question—which, as we argue, our community is well-equipped to answer—holds the key to unlocking unimaginable productivity gains from agents being the primary mechanism we use to interact with data.
Thankfully, while agentic speculation represents a new challenge for data systems, its characteristics present new opportunites for the redesign of data systems. As we show, agentic speculation:
(1) can be high throughput, benefiting from a lot of requests to the backend systems, issued in sequence and/or in parallel, to determine how to solve the given task.
(2) is heterogeneous, spanning coarse-grained data and metadata exploration, partial and complete solution formulation, and validation—allowing LLM agents to make progress with approximate or incomplete outputs in early stages.
(3) has redundancy: many requests may access similar data or perform overlapping operations, offering opportunities to share computation or eliminate redundant work.
(4) is steerable: since speculation is fundamentally exploratory, if we move beyond the query-answer paradigm and allow data systems to more directly communicate with LLM agents, it could help steer LLM requests toward the most promising directions.
In this paper, we propose a new research vision for our community around redesigning data systems for agents, by leveraging the aforementioned characteristics of speculation—scale, heterogeneity, redundancy, and steerability. In Sec. 2, we illustrate through case studies the characteristics of present-day agentic speculation. In Sec. 3, we propose a new architecture for agent-first data systems. In Sec. 4, 5, and 6, we identify new research opportunities in the interface, query processing, and storage layers, respectively.
2. Case Studies
In this section we explore the characteristics of agentic workloads through two case studies—and identify patterns in these queries that present optimization opportunities. While these case studies are simple, they are easier to evaluate for correctness.
(a) Success @ K
(b) Success vs. Turns
Figure 1. Results on the BIRD dataset
We use the BIRD text2SQL benchmark (Li et al, 2023) in our first study. We wanted to explore if present-day LLMs benefit from increasing the number of requests—in parallel or in sequence. We used DuckDB as our backend, and GPT-4o-mini and Qwen2.5-Coder-7B-Instruct as two LLMs. To first evaluate parallel requests, we simulated the behavior of an LLM agent “in charge,” with a number of “field” agents each independently attempting the task, followed by the agent-in-charge picking one among the corresponding solutions. We plot the average success rate versus the number of LLM attempts in Figure 1(a). To instead evaluate sequential requests, we had a single LLM agent issue queries until it was satisfied and once again plot the success rate versus the number of steps taken in Figure 1(b). We find that:
Agentic speculation—in sequence or in parallel—helps improve accuracy.
The success rate of agentic workloads increases as a function of requests, and by 14%–70% in our case study.
(a) versus subexpression size.
(b) versus root operation.
Figure 2. Total vs. unique subexpressions (count and proportion) across 50 attempts generated by GPT-4o-mini per problem, aggregated over the full BIRD dataset. Here, PR=Projection, TS=Scan, FI=Filter, HJ=Hash Join, UA=Aggregate, OT=other operations.
Next, we quantify the degree to which work sharing is possible across requests. We focus our attention on the parallel setting, with 50 independent attempts—and evaluate the redundancy across these attempts. We plot the total number and distinct number of sub-plans or sub-expressions of each size in the 50 query plans generated for a given task, aggregated across the full BIRD dataset, in Figure 2(a). We present a similar plot for sub-plans grouped by root operator type in Figure 2(b). We find:
Agentic speculation has substantial redundancy across requests.
Across queries, the number of distinct sub-plans of each size is often a small fraction of less than 10-20% of the total, representing considerable potential for sharing computation.
Figure 3. Labeled agent activities, with x-axis showing normalized position in the trace, and each row (activity) normalized independently. Agents first explore table and columns then formulate queries, with phases often overlapping.
Our second case study is more involved than text2SQL and helps us study the phases of agentic speculation. We evaluate the performance of a data agent that must combine information from two separate backend databases, chosen from PostgreSQL, SQLite, MongoDB, and DuckDB. For example, one task involves cleaning customer information from MongoDB to join with user interaction data (e.g., upvotes) in DuckDB. As such, it is impossible to complete this task in a single shot, and successful attempts typically involve interacting with both backends, followed by some computation in Python. We collected 44 sequential traces of OpenAI’s o3 model attempting each of the 22 tasks twice, with about half resulting in correct answers. We then manually labeled each action taken by the LLM with an annotation: exploring metadata and sample data (targeting schemas or with LIMIT), exploring column statistics (distinct values or aggregates), attempting part of the query, or all of it. As we can see in the aggregated heatmap of traces in Figure 3, exploring metadata and sample data typically happens first, followed by statistics, after which the next two phases emerge. However, these phases are not clearly delineated, and each phase is present throughout the trace. So we find:
Agentic speculation is heterogeneous in its information needs.
Requests from agents vary greatly in the information necessary, from coarse-grained exploration of metadata and data statistics, to partial or more complete attempts at addressing the task. Coarse-grained, exploratory requests typically happen early on.
In the following, we describe the earlier phases as metadata exploration, and the latter phases as solution formulation.
Table 1. Mean activity counts per agent trace, averaged across all traces, with and without human expert-provided hints.
Activity Avg (No Hints) Avg (w/ Hints) Reduction (%)
exploring tables 3.44 2.95 -14.2
exploring specific columns 3.56 2.57 -27.7
attempting part of the query 4.28 2.71 -36.6
attempting entire query 1.26 1.05 -16.6
all SQL queries 12.67 10.38 -18.1
Next, we wanted to explore if grounding provided by the backend system could help reduce the number of steps taken to reach the solution. So, we simulated this by measuring the impact of injecting hints into the prompt, where the hint provides background information useful for the task, such as which column contains information pertinent to the task. Again, we collected 44 sequential traces (two per task) with hints provided, and then measured the average number of steps required across attempts and tasks when hints were provided versus not. As shown in Table 1, the impact of hints is substantial. We find:
Agentic speculation is steerable through grounding hints.
Speculation traces can become much more efficient—reducing queries by >>20%, depending on phase—if proactively provided grounding pertinent to the task.
Based on the characteristics gleaned via our case studies, we next propose a new architecture for agent-first data systems.
3. Agent-First Data System Architecture
Here, we outline a potential architecture for a data system that is agent-first, as shown in Figure 4.
Given a user task, an army of LLM agents can issue one or more probes to the backend system, possibly associated with relative priorities. We call these probes rather than queries for two reasons. First, they could go beyond SQL in providing background information about the nature of the request, such as the phase (metadata exploration or solution formulation), the identity of the agent issuing the request, the degree of accuracy required, overall goals, among others. We envision this information to be specified in natural language or some other flexible format to be interpreted by in-database agents. Second, the probes could go beyond SQL on data or metadata (e.g., via information_schema) to search for tokens that may be present in any table (either column or row) to help identify which tables need to be accessed.
Then, these probes are parsed and interpreted by an agentic interpreter component within the database. For each of these probes, the system could provide answers, possibly approximate, and also proactively provide information going beyond answers to help steer the agents through grounding feedback. We describe our interface as well as proactive feedback in Sec. 4.
Given one or more probes, our probe optimizer attempts to satisfice, i.e., produce reasonable results that address needs, without evaluating the query completely, as described in Sec. 5; this optimizer leverages and extends traditional database research on multi-query optimization and approximate query processing.
To improve efficiencies, the storage and transactional components of our data systems will need to evolve, as described in Sec. 6. We introduce an agentic memory store to store any grounding gleaned, so that they can be used in future probes. For updates, our shared transaction manager efficiently handles the sheer redundancy in state involved across many potential transactions.
Figure 4. Agent-First Data Systems Architecture; components that are dashed involve LLM agents. Boxes in pink are covered in Sec. 4; blue in Sec. 5; orange in Sec. 6.
4. Query Interfaces
In this section, we focus on agent-database interaction. We start by describing how probes (i.e., input from agents to the data system) need to go beyond SQL in Section 4.1. Then, we discuss how data systems can go beyond the query-result paradigm in providing additional grounding information to help steer the agents in Section 4.2.
4.1. From Agents to Data Systems
Probes from agents need to go beyond SQL in specifying why or how a given query needs to be answered. Moreover, for certain types of information needs, SQL may be limiting, necessitating new operators. We describe each aspect in turn.
Providing Background Information. If all an agent can do is specify SQL queries, then all the data system can do is provide exact results for those queries, making speculative probing inefficient. While specifying LIMIT or exact degree of approximation is one option, it provides limited expressive power. Therefore, as part of a probe, agents can specify one or more SQL queries, along with what we call a brief: natural language statements about the probe’s goals and intents, its current phase (metadata exploration or solution formulation), approximation needs and priorities across queries or probes, as well as any other open-ended information. These briefs are in turn examined by the probe interpreter agent within the data system and used to guide optimization and execution, e.g., what order to execute the queries (if at all) and degree of approximation (or accuracy) based on goals and phase. Determining how to set accuracy based on this natural language input is an open question and needs to also take into account relative query execution costs.
Across a batch of queries specified within a probe, the probe can additionally specify open-ended goals that go beyond simple accuracy, such as pair-wise priorities, or indicating that only $k$ query among $n$ specific queries needs to be performed to completion (and the data system can decide which one to maximize efficiency). For example, if a field agent in an exploratory phase wants to get a sense for the differences in sales performance of stores on the US West coast vs. East coast, it can specify, as a part of the probe, that the data system needs to generate statistics for two states each from each coast, with the data system being able to pick which ones. The interface can furthermore allow for other forms of approximation indicators that are time-consuming for humans to write but can now be done by agents, e.g., specifying termination criteria, functions that the data system can evaluate on the partial result sets to know if some queries can be terminated early. For example, one termination criteria could be defined to stop execution of multiple “needle-in-a-haystack” type queries mid-execution because the answers are too similar to previous ones (where an agent defined function is evaluated on partial result sets to determine answer similarities).
Extending Capabilities through Flexible Probes. In many cases, agents are unsure of even where to start and which tables to query for a given task—because they lack knowledge of how the data is organized. Suppose an agent is tasked with finding out how a given company will be “impacted by increased tariffs on the import of electronic goods.” This agent may want to find tables whose name is semantically similar to “electronics,” or whose rows contain data that is semantically similar. Such probes that ask for semantically similar contents—be it tables, columns, or rows—to a specific phrase, located anywhere, are impossible to address within SQL, but are valuable during the early exploratory phases. Thus, we need native support for semantic similarity operators, beyond LIKE, where the operators are applied to any data or metadata in the data system. Furthermore, as we will discuss in Sec. 6, the agents will rely on metadata stored in agentic memory, on cells, rows, columns, and tables, typically written by agents themselves—to understand data semantics, and as such will need to frequently query or update this metadata. Although the above functionalities may be possible through a combination of tools (e.g., store metadata separately in a vector database, look it up and then issue SQL queries), determining what and how to actually store, and how to keep it up-to-date is a challenge. Moreover, a data system that holistically supports all data and metadata needs can be more effectively used by agents.
4.2. From Data Systems to Agents
In addition to simply answering probes, data systems should steer agents towards better probes, which in turn can lead to improved efficiencies. In this way, the data system acts in a more proactive (Zeighami et al, 2025) manner, akin to how a data engineer or administrator may assist data analysts in satisfying information needs as efficiently as possible. This information can be provided in addition to, or in lieu of the answers to the probes, in natural language. This steering can serve two purposes: (1) helping agents by providing auxiliary data-centric information the data system finds relevant, as a side-channel, and (2) providing feedback to agents on efficiency and costs to assist the agents in designing their probes. We envision sleeper agents within the data system that are invoked on-demand to gather information in parallel with answering probes, to be returned in addition to probe answers, as we discuss below.
Auxiliary Information. As we saw in Table 1, providing grounding hints or feedback can reduce the number of probes agents need to complete their tasks. We envision sleeper agents tasked with identifying and providing such hints as auxiliary information along with answers. For example, the sleeper agent could find and share other related tables—to be either joined with (as in join discovery, e.g., (Sarma et al, 2012)), or replacing the current table as the focus of analysis, especially if the current table proves irrelevant. Or rather than the agent having to guess why they got an empty result, the sleeper agent can provide feedback reminiscent of why-not provenance (Cheney et al, 2009), e.g., the probe assumed that states were encoded with two letter acronyms like “CA”, but instead they are listed out in entirety.
Cost Estimates and Cost-Based Feedback. Grounding can also come in the form of cost estimates; for example, even before executing a query, estimated costs (especially if higher than expected) can be provided to the agent to help determine if the probe must be run to completion, and suggest the agent to modify the probe (e.g., to just focus on California instead of all of USA), or increase the degree of approximation. This can similarly be applied across probes. For example, if the sleeper agent predicts that the probes are performing a set of tasks in sequence, it can suggest to the field agent to batch them, if it proves to be cheaper. The sleeper agent can also take into account related materialized answers, or if a similar query was just answered for another agent. In such cases, the sleeper agent can suggest modifying the input probe to probes with such pre-defined answers to improve efficiency—or it can output the answer for such related probes in the side-channel.
Next, we discuss how to efficiently provide answers to probes.
5. Processing and Optimizing Probes
As discussed in Sec. 2, agentic probes will have much higher throughput than those issued by human sources (e.g., web applications). Importantly, in agent-first data systems, our goal is not to optimize overall throughput as in traditional databases, but to evaluate probes enough such that agents can make their decision on how to proceed in the next turn. With that in mind, this section discusses what needs to change in data systems to effectively support probes.
5.1. Supporting Exploration
Our agentic probes will consist of exploratory queries to establish grounding. Some explorations will inevitably be cast in natural language (NL) as agents lack knowledge about the underlying databases (e.g., “how to find out how many tables are stored?”) with others expressed using SQL (e.g., SELECT count(*) FROM
information_schema.tables
in PostgreSQL). Today’s databases are not designed to answer NL queries. The probe optimizer in our agent-first data system must therefore orchestrate the mix of NL and SQL queries by utilizing different agents at the scale of probes.
To illustrate, consider identifying the stores that show an increasing sales trend. Our agents will first need to find out which tables are used to store sales data. A straw person probe execution plan is to pose NL questions to a web search agent to discover how to look up table schema, and execute the found queries on our database. While these are simple queries on our database’s metadata tables, the outputs returned from such queries often contain lots of unnecessary information. For instance, PostgreSQL maintains hundreds of internal tables and indexes even without any user table defined. Coupled with the user tables, the results can easily grow to thousands of rows. Feeding all the rows to our query formulation agent is simply a waste of its limited context length.
Subsequently, to discover what constitutes an increasing trend, one strategy is to find examples of “trend queries” using NL with a web search agent, then feed the returned information to a query formulation agent to translate into SQL. We will likely get lots of example queries online, and our database will be bombarded with lots of inapplicable queries (e.g., they refer to non-existent tables, or identify the wrong trend). Worse yet, all such explorations will be mixed with other agents formulating solutions. With today’s data systems, we have no means to identify which queries are part of agentic exploration (and hence do not need to be evaluated completely). We envision that our probe optimizer will prioritize queries based on their phases (i.e., a form of admission control). Furthermore, we will store previously gleaned information using our agentic memory store to avoid repeated querying of the same information, and train agents to query our memory store instead of including such information as part of the prompt each time.
5.2. Probe Optimization
As mentioned, probes issued by agents, unlike queries issued by humans, do not require complete answers. The database interface allows the agents to specify goals, and approximation needs in natural language via briefs, which are then used by the database to decide which probes to execute and to what degree of accuracy. This means the goal of the query optimizer, unlike in traditional data systems, is to decide both what queries to execute (and to what degree of approximation) to satisfice for the probe, as well as how to execute them. In doing so, the optimization has a new objective: minimize the total time spent on answering the field agents’ probes given available computational resources. Solving this optimization problem requires the database internally balancing cost/accuracy trade-offs: if the database chooses to answer a query with high degree of approximation providing insufficient answers to save cost upfront, the agent may ask many follow-ups with increased accuracy requirements, thus increasing total time spent answering the agent’s probes. We next discuss how we envision such an optimization problem can be solved, within a given batch of probes sampled at an interaction turn (Section 5.2.1), and across batches of probes across turns and agents (Section 5.2.2).
5.2.1. Intra-Probe Optimization
We first discuss how to optimize a given batch of probes to provide sufficient information for the agent while minimizing computational cost.
Deciding on What to Execute. The database must first decide what queries to run and to what degree of approximation, taking the probe and its briefs into account. This requires the database to reason about the data and probe semantics, including the agent’s goals and phases. To do so, the database can use semantic query and data understanding to check if they match user’s intents, and prune away queries it deems not semantically meaningful. For instance, during the exploration phase, the database can examine the projected columns in probe to see if they are relevant to the user’s intent, and if not prune such columns, or the probes away as a whole. Moreover, the database can compare probes within a batch, guided by probe’s briefs that may have specified approximation needs across probes. The database can then make cost estimates and compare information gain from the probes to decide which probes are more helpful and/or cheaper. For example, given two probes $P$ and $P^{\prime}$ the database can prune $P^{\prime}$ away if rows that would be returned by $P^{\prime}-P$ are deemed irrelevant to the agent’s goal. This is reminiscent of prior work on pruning queries as part of visualization recommendation (et al, 2015), and deciding query equivalence as part of query synthesis given user provided input/output examples (Zloof, 1975; Wang et al, 2017), although the scale of queries to compute the differences will be much larger in agentic workloads. Finally, the database can take the agent’s phase into account; for example, return coarse grain approximations during exploration, but more accurate answers during solution formulation. Beyond pruning queries, we envision agents will be able to examine other internal database states (e.g., buffer pool, outputs of query operators) to determine if it should continue with query evaluation, or move on to the next turn.
Efficient Execution. As mentioned in Sec. 2, probes have substantial redundancy that we can exploit by sharing computation across them. Multi-query optimization (et al, 2012; Sellis, 1988), approximate query processing (Garofalakis et al, 2001) and caching of partial query results can be used to improve efficiency. However, there are new unique challenges. For example, different probes will have different approximation requirements and may be accompanied with termination criteria (a function that can be evaluated on partial results to know if they are sufficient, see Sec. 4), which makes it more difficult to reason about their semantics and what can be shared. Moreover, the database can incrementally evaluate queries, reminiscent of incremental query processing (Chandramouli et al, 2014), but with the new challenge of decision making across them; e.g., the database must decide which probe is the most useful to the agent and provide higher accuracy for that probe first before increasing accuracy for other probes. Finally, query planning and processing can be done jointly with optimization, e.g., the database can re-evaluate its decisions on what queries to run, or increase its level of approximation for some query during planning or processing as it obtains more information.
5.2.2. Inter-Probe Optimizations
The database can furthermore leverage the sequential interactions with agents across turns to further optimize both the queries it decides to run and their execution.
Deciding on What to Execute. Besides the strategies discussed in Section 5.2.1, the database can consider all interactions with the agent to decide what queries to run. First, it can decide on queries to run based on whether they provide any new information given past queries answered. For example, when given probes $P$ and $P^{\prime}$ by the agents across consecutive turns, if the output between $P$ and $P^{\prime}$ is expected to not provide new information to the agent—e.g., $P^{\prime}$ adds new columns that are non-relevant to the agent’s goal—then $P^{\prime}$ can be dropped. Furthermore, the database can decide what queries to run in order to minimize the number of future follow-up probes. For example, based on the agent’s goal specified in the probe briefs, it can run a query it finds maximally useful to the agent exactly and to completion rather than approximately even if the current query may take longer, expecting that the extra computation upfront will reduce total runtime across future interactions with the agent.
Efficient Execution. The database can decide to materialize and cache answers by observing the query history and considering the agent’s intent. For example, based on the history and the agent’s intent, the database can expect future probes will continue to involve the join of certain tables and can materialize the join.
6. Indexing, Storage, and Transactions
The heterogeneity and redundancy of agentic speculation workloads fundamentally challenge the assumptions of the storage layer of today’s data stores, specifically, that workloads are static and independent.
For static workloads, data systems rely on predefined indexes and fixed storage layouts (e.g., column-based for OLAP) based on recurring workload patterns. Agentic probes, by contrast, evolve from coarse-grained metadata exploration to final validation. This dynamism makes static tuning ineffective. Meanwhile, the exploratory (resp. solution formulation) phases of different probes may be similar and can benefit from similar layouts.
On the independence front, data systems treat queries as unrelated, such that concurrent access (specifically writes) from these queries must be isolated from each other. While this simplifies application logic and ensures consistency, these mechanisms prevent cooperative sharing of state with rare exceptions (Gupta et al, 2011). Instead of isolation, agentic workloads demand a more cooperative model—one that can safely share intermediate state across different probes, many of which are likely to be similar.
Hence, we propose two key ideas to improve performance. First, we propose an agentic memory store that acts as a “pseudo-index” to help agentic probes quickly find information that may be helpful, either directly accessed by them, or on their behalf by sleeper agents. Second, we propose a new transactions framework that is centered on state sharing across probes, each of which may be independently attempting to complete a user-defined sequence of updates.
6.1. Agentic Memory Store
The exploratory phase of agentic speculation aims to identify the right tables and columns to operate on. To improve efficiency, data systems should maintain a persistent, queryable agentic memory store—a semantic cache that provides grounding.
Artifacts. The first question is what should be stored. One idea is to store the results of prior probes and partial solutions, so that agents can reuse what is known about the data and metadata, enabling similar probes to be more efficient. In addition, we can store information about the data and metadata, possibly associated with the tables themselves. We can store encoding formats for columns, missing value information, and time and location granularities. For example, an agent trying to explore various sales partitions may retrieve a number of them, along with the metadata in the agentic memory that indicates the date ranges or location ranges associated with each—so that it can make a more informed decision about which ones to probe further.
To implement this store, we can embed the agentic metadata with the table directly, to be retrieved if the table is queried. For all other open-ended information, one approach is to use a vector index to support semantic similarity search on embeddings (e.g., querying with a probe might retrieve other similar probes, and what worked for them). However, this approach may not work as well for more targeted or more structured lookups.
Updates to the Store. A separate concern is how this memory store is maintained during updates. Updates could be in the form of new probes being executed that may provide new information that augments or supersedes existing ones. Or, it could be to the underlying data or metadata, necessitating updates to any related information in the agentic memory. For example, if there is a schema update, the results of a prior probe that used that table may no longer be relevant. One approach is to allow this memory store be inconsistent with the data/metadata, and instead be updated by any new probes that discover that the information is stale. However, the downside is that the stale information may lead a new probe to make a mistake. For example, suppose the agentic memory indicates that the only relevant sales information can be found in three tables, but after that, additional relevant tables were added; here, new probes may end up returning incorrect results. Additional challenges emerge in supporting access control for multiple users. For example, agents acting on different users’ behalf could ask similar questions (e.g., ”Where is the employee’s availability stored?”). Sharing answers across such agents boosts efficiency—but raises privacy concerns, especially in the aggregate (Popa et al, 2011). Addressing these challenges will need to draw inspiration from work on knowledge bases as well as schema evolution.
6.2. Performing Branched Updates
[…]
Branch Isolation. Existing models of branching consistency, developed in the weak consistency era, e.g., in Bayou, Dynamo, or Tardis (Petersen et al, 1996; DeCandia et al, 2007; Crooks et al, 2016), can offer inspiration. However, agentic speculation goes further: multiple agents may create forks that must eventually reconcile—not just with the mainline, but with each other. This requires new models of multi-agent, multi-version isolation. Most branches will be similar—e.g., same schema, 90% identical data—but isolation requires that their effects remain logically separate.
Efficient forking and rollbacks. Naively duplicating entire databases per branch is prohibitively expensive and inefficient, making support for efficient forking crucial. Industrial systems like Neon (Neon, [n. d.]), Aurora (Verbitski et al, 2017) and academic systems like Tardis (Crooks et al, 2016) adopt copy-on-write approaches to lazily clone state. However, these are still far from what is needed for massive speculation. We need new concurrency mechanisms that exploit similarity across branches and preserve logical isolation (no cross-contamination), to enable massive parallel forking. This is analogous to MVCC on steroids: forking possibly thousands of near-identical snapshots and rolling back all but one. Unlike traditional data systems, where rollbacks are rare, we require ultra-fast rollbacks (i.e., fast aborts for failed branches).
7. Conclusion
We described our vision for a data system to support emerging agentic workloads. These workloads involve massive speculative probing, characterized by a high-throughput redundant mix of discovery and validation, specified by a combination of queries and natural language. We present an architecture for such redesigned data systems, and discuss emergent research challenges.

### v2

@@
-Serving Our AI Overlords: Redesigning Data Systems for Agents
+Supporting Our AI Overlords: Redesigning Data Systems to be Agent-First
 Abstract.
@@
 Figure 1. Results on the BIRD dataset
-We use the BIRD text2SQL benchmark (Li et al, 2023) in our first study.
+We use the BIRD text2SQL benchmark (bird) in our first study.
 We wanted to explore if present-day LLMs benefit from increasing the number of requests—in parallel or in sequence.
@@
 In addition to simply answering probes, data systems should steer agents towards better probes, which in turn can lead to improved efficiencies.
-In this way, the data system acts in a more proactive (Zeighami et al, 2025) manner, akin to how a data engineer or administrator may assist data analysts in satisfying information needs as efficiently as possible.
+In this way, the data system acts in a more proactive (zeighami2025llm) manner, akin to how a data engineer or administrator may assist data analysts in satisfying information needs as efficiently as possible.
 This information can be provided in addition to, or in lieu of the answers to the probes, in natural language.
@@
 We envision sleeper agents tasked with identifying and providing such hints as auxiliary information along with answers.
-For example, the sleeper agent could find and share other related tables—to be either joined with (as in join discovery, e.g., (Sarma et al, 2012)), or replacing the current table as the focus of analysis, especially if the current table proves irrelevant.
-Or rather than the agent having to guess why they got an empty result, the sleeper agent can provide feedback reminiscent of why-not provenance (Cheney et al, 2009), e.g., the probe assumed that states were encoded with two letter acronyms like “CA”, but instead they are listed out in entirety.
+For example, the sleeper agent could find and share other related tables—to be either joined with (as in join discovery, e.g., (sarma2012finding)), or replacing the current table as the focus of analysis, especially if the current table proves irrelevant.
+Or rather than the agent having to guess why they got an empty result, the sleeper agent can provide feedback reminiscent of why-not provenance (cheney2009provenance), e.g., the probe assumed that states were encoded with two letter acronyms like “CA”, but instead they are listed out in entirety.
 Cost Estimates and Cost-Based Feedback.
@@
 Our agentic probes will consist of exploratory queries to establish grounding.
-Some explorations will inevitably be cast in natural language (NL) as agents lack knowledge about the underlying databases (e.g., “how to find out how many tables are stored?”) with others expressed using SQL (e.g., SELECT count(*) FROM
-information_schema.tables
-in PostgreSQL).
+Some explorations will inevitably be cast in natural language (NL) as agents may lack knowledge about the underlying databases (e.g., “how to find out how many tables are stored?”) with others expressed using SQL (e.g., SELECT count(*) FROM information_schema.tables in PostgreSQL).
 Today’s databases are not designed to answer NL queries.
@@
 Our agents will first need to find out which tables are used to store sales data.
-A straw person probe execution plan is to pose NL questions to a web search agent to discover how to look up table schema, and execute the found queries on our database.
+A straw person probe execution plan is to pose NL questions to a web search agent to discover how to look up table schema for our specific database dialect, and execute the found queries on our database.
 While these are simple queries on our database’s metadata tables, the outputs returned from such queries often contain lots of unnecessary information.
 For instance, PostgreSQL maintains hundreds of internal tables and indexes even without any user table defined.
-Coupled with the user tables, the results can easily grow to thousands of rows.
-Feeding all the rows to our query formulation agent is simply a waste of its limited context length.
-Subsequently, to discover what constitutes an increasing trend, one strategy is to find examples of “trend queries” using NL with a web search agent, then feed the returned information to a query formulation agent to translate into SQL.
+Coupled with the user tables, the results can easily grow to thousands—or hundreds of thousands—of rows.
+Feeding all the rows to our query formulation agent is a waste of its limited context length.
+As mentioned in Section 4.1, we further need the ability to query tokens regardless of where they appear across databases, be it as part of metadata or data.
+Subsequently, to discover what constitutes an increasing trend, one strategy is to find examples of “trend queries” (possibly using window queries) using NL with a web search agent, then feed the returned information to a query formulation agent to translate into SQL.
 We will likely get lots of example queries online, and our database will be bombarded with lots of inapplicable queries (e.g., they refer to non-existent tables, or identify the wrong trend).
@@
 We first discuss how to optimize a given batch of probes to provide sufficient information for the agent while minimizing computational cost.
-Deciding on What to Execute.
+Deciding What to Execute.
 The database must first decide what queries to run and to what degree of approximation, taking the probe and its briefs into account.
@@
 For example, given two probes $P$ and $P^{\prime}$ the database can prune $P^{\prime}$ away if rows that would be returned by $P^{\prime}-P$ are deemed irrelevant to the agent’s goal.
-This is reminiscent of prior work on pruning queries as part of visualization recommendation (et al, 2015), and deciding query equivalence as part of query synthesis given user provided input/output examples (Zloof, 1975; Wang et al, 2017), although the scale of queries to compute the differences will be much larger in agentic workloads.
+This is reminiscent of prior work on pruning queries as part of visualization recommendation (seedb), and deciding query equivalence as part of query synthesis given user provided input/output examples (zloof; scythe1), although the scale of queries to compute the differences will be much larger in agentic workloads.
 Finally, the database can take the agent’s phase into account; for example, return coarse grain approximations during exploration, but more accurate answers during solution formulation.
@@
 2, probes have substantial redundancy that we can exploit by sharing computation across them.
-Multi-query optimization (et al, 2012; Sellis, 1988), approximate query processing (Garofalakis et al, 2001) and caching of partial query results can be used to improve efficiency.
+Multi-query optimization (sharedb; mqo1; mqo2), approximate query processing (aqp) and caching of partial query results can be used to improve efficiency.
 However, there are new unique challenges.
@@
 4), which makes it more difficult to reason about their semantics and what can be shared.
-Moreover, the database can incrementally evaluate queries, reminiscent of incremental query processing (Chandramouli et al, 2014), but with the new challenge of decision making across them; e.g., the database must decide which probe is the most useful to the agent and provide higher accuracy for that probe first before increasing accuracy for other probes.
+Moreover, the database can incrementally evaluate queries, reminiscent of incremental query processing (trill), but with the new challenge of decision making across them; e.g., the database must decide which probe is the most useful to the agent and provide higher accuracy for that probe first before increasing accuracy for other probes.
 Finally, query planning and processing can be done jointly with optimization, e.g., the database can re-evaluate its decisions on what queries to run, or increase its level of approximation for some query during planning or processing as it obtains more information.
@@
 The database can furthermore leverage the sequential interactions with agents across turns to further optimize both the queries it decides to run and their execution.
-Deciding on What to Execute.
+Deciding What to Execute.
 Besides the strategies discussed in Section 5.2.1, the database can consider all interactions with the agent to decide what queries to run.
@@
 For example, based on the agent’s goal specified in the probe briefs, it can run a query it finds maximally useful to the agent exactly and to completion rather than approximately even if the current query may take longer, expecting that the extra computation upfront will reduce total runtime across future interactions with the agent.
+Yet another direction is to treat the problem as one of exploration vs.
+exploitation: instead of always trying to provide rapid answers to queries by satisficing, the database can sometimes prioritize exploration of underexplored solution spaces to identify those solutions that have an unanticipated benefit, in order to maximize utility over time.
 Efficient Execution.
@@
 On the independence front, data systems treat queries as unrelated, such that concurrent access (specifically writes) from these queries must be isolated from each other.
-While this simplifies application logic and ensures consistency, these mechanisms prevent cooperative sharing of state with rare exceptions (Gupta et al, 2011).
+While this simplifies application logic and ensures consistency, these mechanisms prevent cooperative sharing of state with rare exceptions (entangled).
 Instead of isolation, agentic workloads demand a more cooperative model—one that can safely share intermediate state across different probes, many of which are likely to be similar.
@@
 For example, agents acting on different users’ behalf could ask similar questions (e.g., ”Where is the employee’s availability stored?”).
-Sharing answers across such agents boosts efficiency—but raises privacy concerns, especially in the aggregate (Popa et al, 2011).
+Sharing answers across such agents boosts efficiency—but raises privacy concerns, especially in the aggregate (cryptdb).
 Addressing these challenges will need to draw inspiration from work on knowledge bases as well as schema evolution.
@@
 Branch Isolation.
-Existing models of branching consistency, developed in the weak consistency era, e.g., in Bayou, Dynamo, or Tardis (Petersen et al, 1996; DeCandia et al, 2007; Crooks et al, 2016), can offer inspiration.
+Existing models of branching consistency, developed in the weak consistency era, e.g., in Bayou, Dynamo, or Tardis (bayou; dynamo; tardis), as well as versioned databases (huang2017orpheusdb) can offer inspiration.
 However, agentic speculation goes further: multiple agents may create forks that must eventually reconcile—not just with the mainline, but with each other.
@@
 Naively duplicating entire databases per branch is prohibitively expensive and inefficient, making support for efficient forking crucial.
-Industrial systems like Neon (Neon, [n.
-d.]), Aurora (Verbitski et al, 2017) and academic systems like Tardis (Crooks et al, 2016) adopt copy-on-write approaches to lazily clone state.
-However, these are still far from what is needed for massive speculation.
+Industrial systems like Neon (neon), Aurora (aurora), and Bauplan (tagliabue2025safe) and academic systems like Tardis (tardis) adopt copy-on-write approaches to lazily clone state.
+However, these are still far from what is needed for agentic speculation at a massive scale.
 We need new concurrency mechanisms that exploit similarity across branches and preserve logical isolation (no cross-contamination), to enable massive parallel forking.
@@
 7. Conclusion
-We described our vision for a data system to support emerging agentic workloads.
-These workloads involve massive speculative probing, characterized by a high-throughput redundant mix of discovery and validation, specified by a combination of queries and natural language.
-We present an architecture for such redesigned data systems, and discuss emergent research challenges.
+We described our vision for data systems that natively support emerging agentic workloads.
+These workloads involve agentic speculation, characterized by a high-throughput, heterogeneous, and redundant mix of discovery and validation, specified by probes ideally involving a combination of queries and natural language.
+We present one such architecture for such redesigned data systems, and discuss emergent research challenges.
+Acknowledgments
+This work is supported by NSF grants IIS-1955488, IIS-2027575, DGE-2243822, IIS-2129008, IIS-1940759, and IIS-1940757, DOE awards DE-SC0016260, AC02-05CH11231, DARPA agreement HR00112590131, and funds from the state of California.
+This work is also supported by EPIC Data Lab sponsors and affiliates, including Adobe, Bridgewater, Google, G-Research, Microsoft, PromptQL, Sigma Computing, and Snowflake, as well as Berkeley Sky Lab sponsors and affiliates, including Accenture, AMD, Anyscale, Broadcom, Google, IBM, Intel, Intesa Sanpaolo, Lambda, Lightspeed, Mibura, Samsung SDS, SAP, Cisco, Microsoft and NVIDIA.
+Compute credits were provided by Azure, Modal, NSF (via NAIRR), and OpenAI.
+We thank the reviewers for their feedback, as well as Arash Nourian for helpful discussions.